OpenStudy (anonymous):

alright tricky one! Find f(6) if f(x) = x^2 ÷ 3 + x A.4 B.10 C.18

OpenStudy (anonymous):

how do you solve it

OpenStudy (goformit100):

sorry i went wrong

OpenStudy (jackellyn):

=(6)^2 / 3 + 6 =36 /3 +6 =12 +6 = 18

OpenStudy (anonymous):